# Resolving the QCD Axion Domain Wall Problem with a Light Axion

### v1

@@ IV.1 Setup
where $n_{a}$ and $n_{\phi}$ are integers. Here, $V_{\mathrm{QCD}}$ and $V_{\Lambda}$ come from the non-perturbative effects of QCD and the hidden sector, respectively. We set the origins of $a$ and $\phi$ so that $a=\phi=0$ corresponds to the potential minimum. For later convenience, we define
@@
$=\frac{1}{2}\begin{pmatrix}a&\phi\end{pmatrix}\mathcal{M}^{2}\begin{pmatrix}a\\ \phi\end{pmatrix}\ ,$ (25)
@@
$\mathcal{M}^{2}=\begin{pmatrix}m_{a}^{2}+\dfrac{n_{a}^{2}\Lambda^{4}}{f_{a}^{2}}&\dfrac{n_{a}n_{\phi}\Lambda^{4}}{f_{a}f_{\phi}}\vspace{2mm}\\ \dfrac{n_{a}n_{\phi}\Lambda^{4}}{f_{a}f_{\phi}}&\dfrac{n_{\phi}^{2}\Lambda^{4}}{f_{\phi}^{2}}\end{pmatrix}=\begin{pmatrix}m_{a}^{2}+m_{a,\Lambda}^{2}&\epsilon m_{a,\Lambda}^{2}\vspace{2mm}\\ \epsilon m_{a,\Lambda}^{2}&\epsilon^{2}m_{a,\Lambda}^{2}\end{pmatrix}\ .$ (26)
@@
$M_{d}^{2}\equiv\begin{pmatrix}m_{H}^{2}&0\vspace{2mm}\\ 0&m_{L}^{2}\end{pmatrix}\equiv\mathcal{O}\mathcal{M}^{2}\mathcal{O}^{T}\ ,$ (27)
@@
$\begin{pmatrix}\phi_{H}\\ \phi_{L}\end{pmatrix}=\mathcal{O}\begin{pmatrix}a\\ \phi\end{pmatrix}\simeq\begin{pmatrix}a+\epsilon\frac{m_{a,\Lambda}^{2}}{m_{a}^{2}+m_{a,\Lambda}^{2}}\phi\\ \phi-\epsilon\frac{m_{a,\Lambda}^{2}}{m_{a}^{2}+m_{a,\Lambda}^{2}}a\end{pmatrix}\ .$ (32)
@@ IV.2 Dynamics of the string-wall network
Once $\phi$ begins to evolve due to $V_{\Lambda}$, both $a$ and $\phi$ roll down to the degenerate minima within each domain, and no potential bias remains. In contrast, while $\phi$ is still frozen at its initial value due to its small mass, the potential energy generally varies across different domains. This energy difference acts as a (transient) potential bias, which can induce the collapse of the string-wall network. In this sense, the setup effectively introduces a transient explicit breaking of the PQ symmetry in the early Universe, which disappears later due to the dynamics of the lighter ALP.
@@
$N_{\rm DW}=2,\quad n_{a}=1,\quad n_{\phi}=1\ .$ (33)
We further assume that the lighter degree of freedom remains frozen until well after the domain wall network collapses.444 In fact, this is a rather conservative condition for the collapse of domain walls. The transient bias could induce a population bias in the domain wall network, which results in a subsequent decay of the domain walls. The decay of domain walls due to population bias for various types of initial fluctuations was studied in Refs. [13, 67]. The onset of $\phi$ oscillation is estimated by
@@
When domain walls are formed with their tension determined by $V_{\mathrm{QCD}}$, the field $a$ takes values $a\simeq 0$ or $\pi f_{a}$ in the two distinct domains, neglecting the contribution from $V_{\Lambda}$. The domain wall tension is given by
@@
Therefore, the string-wall network decays before the transient bias disappears, if $\epsilon$ is sufficiently small.
We then obtain
@@
where we used $m_{L}\simeq\Lambda^{2}/f_{\phi}$ coming from $m_{a,\Lambda}\ll m_{a}$. This inequality, together with the assumption $m_{a}>H_{\rm dec}>m_{L}$, can be satisfied when $f_{\phi}\gg f_{a}$.
@@
Demanding that this accounts for the observed dark matter abundance gives
@@
For the QCD axion with a relatively low decay constant ($f_{a}<10^{10}$ GeV) to be the dominant dark matter, we must have $T_{\rm dec}<T_{\rm QCD}$. We do not consider a lighter QCD axion ($f_{a}>10^{10}$ GeV) for simplicity, where string network and misalignment contributions should also be taken into account.
At the time of collapse, the axion mass is given approximately by its vacuum value:
$m_{a}\approx 2\times\frac{(75.6\,{\rm MeV})^{2}}{f_{a}}\ .$ (43)
@@
The condition $m_{L}<H_{\rm dec}$ requires
$f_{\phi}>f_{\phi,\,\rm min}\sim 2\times 10^{15}\,{\rm GeV}\frac{n_{\phi}}{\sqrt{c}}\left(\frac{1\,{\rm MeV}}{T_{\rm dec}}\right)\left(\frac{10.6}{g_{*,\,\rm dec}}\right)^{1/4}\left(\frac{f_{a}}{10^{8}\,{\rm GeV}}\right)^{1/2}\ .$ (45)
The misalignment mechanism also produces the lighter ALP, with an abundance given by
@@
Avoiding overproduction for a typical initial condition $\phi_{\mathrm{ini}}\sim f_{\phi}$ imposes an upper bound on $f_{\phi}$, defined as $f_{\phi,\,\rm max}$, which corresponds to the largest value of $f_{\phi}$ that yields $\Omega_{L}h^{2}=0.12$. Since $f_{\phi,\,\rm min}$ and $f_{\phi,\,\rm max}$ scale as $1/T^{1/2}_{\rm dec}$ (note that we assume $a$ being the dominant dark matter, which leads to the scaling of $\Lambda^{4}\propto T_{\rm dec}^{3}$ from Eq. (41) with $\sigma H_{\mathrm{dec}}\propto\Lambda^{4}$ and $s_{\mathrm{tot,dec}}\propto T_{\mathrm{dec}}^{3}$), the allowed parameter space is small if the QCD axion is to explain the dark matter. In that case, the light ALP also tends to be dominant, unless the initial value $\phi_{\rm ini}$ is fine-tuned, i.e., $\phi_{\rm ini}\ll f_{\phi}$.
With a mild tuning of $\phi_{\rm ini}/f_{\phi}$, one can suppress $\Omega_{L}$ while keeping $m_{L}$ sufficiently small for the transient bias to be effective. In fact, even when $\phi_{\rm ini}=0$, the transient bias remains nonzero (with $c=1$), as can be seen in Eq. (37).
Alternatively, the light ALP $\phi_{L}$ may account for dark matter while the QCD axion is a subdominant component. This corresponds to taking $f_{a}$ larger than the value in Eq. (42). In this case, choosing $m_{L}\sim H_{\rm dec}$ suppresses the ALP abundance, but $f_{\phi}$ can be increased (which decreases $m_{L}\propto 1/f_{\phi}$) to enhance the ALP abundance ($\propto m_{L}^{1/2}\phi_{\rm ini}^{2}\sim m_{L}^{1/2}f_{\phi}^{2}$) without a tuning for $\phi_{\rm ini}$ and achieve the observed dark matter density. This discussion implies that there should also be the parameter region where both axions become coexisting dominant dark matter.
@@ V Summary and discussion
In this paper we have proposed two simple solutions to the domain-wall problem induced by the QCD axion by introducing a second, massless (or extremely light) axion:
@@
    String-bundle confinement: If the new axion’s PQ symmetry breaks after inflation, the associated strings can be connected to those of the QCD axion by domain walls. The domain walls then terminate on (and are confined by) these strings, forming cosmologically benign “string bundles,” instead of extended domain walls. These bundles behave as cosmic strings and are much less problematic than domain walls, thereby solving the domain wall problem. Such string bundles can survive to the present epoch and generate cosmic birefringence. If the decay constant $f_{\phi}$ is sufficiently large, the string bundles may also leave observable imprints on the CMB anisotropies, providing an additional avenue for detection. A simultaneous detection of the QCD axion and a birefringence signal would be a striking experimental signature.
@@
    Temporary bias mechanism: If the new axion starts with a homogeneous initial value and is sufficiently light, its potential temporarily biases the QCD-axion potential before the light axion begins to oscillate. This induced bias destabilizes and eliminates the walls while leaving the PQ solution to the strong-CP problem intact. In this case, both the QCD axion and ALP can contribute to dark matter depending on the model parameters.
@@
We can also consider the transient bias in a setup where both of the two axions have cosmic strings and string bundles are not effectively formed due to the domain wall numbers larger than unity. In this case, the light axion becomes highly inhomogeneous, and the bias potential is also inhomogeneous even in a single domain. Thus, it is unclear whether the potential bias can still trigger the collapse of the network. To clarify the dynamics of such systems, lattice simulations will be necessary, which is beyond the scope of this work.
@@ Appendix A Typical size of string bundles
as we analytically expected in Sec. III. In the simulation, we choose $f_{\phi}=f_{a}$ to synchronize the string formations of both axions and reduce the computational cost.
Figure 5: The snapshot of the string-wall system that appears in the $N_{\rm DW}=2$ case at an early time (left) and a late time (right). The blue(red) triangles indicate the $a$-($\phi$-)string locations without distinction of the direction, and the black solid lines indicate the domain wall locations. For simplicity, two kinds of axions have the same decay constant, and the mass is constant in the simulations. One can observe that the string bundles are formed in the late time snapshot.

### v2

@@ IV.1 Setup
where $n_{a}$ and $n_{\phi}$ are integers. [^9] The mixing in the coupling to gluons, together with the potential of $\phi$, can be cast into this form by the unitary transformation [36]. Here, $V_{\mathrm{QCD}}$ and $V_{\Lambda}$ come from the non-perturbative effects of QCD and the hidden sector, respectively. We set the origins of $a$ and $\phi$ so that $a=\phi=0$ corresponds to the potential minimum. For later convenience, we define
@@
$=\frac{1}{2}\begin{pmatrix}a&\phi\end{pmatrix}\mathcal{M}^{2}\begin{pmatrix}a\\ \phi\end{pmatrix}\$, (25)
@@
$\mathcal{M}^{2}=\begin{pmatrix}m_{a}^{2}+\dfrac{n_{a}^{2}\Lambda^{4}}{f_{a}^{2}}&\dfrac{n_{a}n_{\phi}\Lambda^{4}}{f_{a}f_{\phi}}\vskip 5.69054pt\\ \dfrac{n_{a}n_{\phi}\Lambda^{4}}{f_{a}f_{\phi}}&\dfrac{n_{\phi}^{2}\Lambda^{4}}{f_{\phi}^{2}}\end{pmatrix}=\begin{pmatrix}m_{a}^{2}+m_{a,\Lambda}^{2}&\epsilon m_{a,\Lambda}^{2}\vskip 5.69054pt\\ \epsilon m_{a,\Lambda}^{2}&\epsilon^{2}m_{a,\Lambda}^{2}\end{pmatrix}\$. (26)
@@
$M_{d}^{2}\equiv\begin{pmatrix}m_{H}^{2}&0\vskip 5.69054pt\\ 0&m_{L}^{2}\end{pmatrix}\equiv\mathcal{O}\mathcal{M}^{2}\mathcal{O}^{T}\ ,$ (27)
@@
$\begin{pmatrix}\phi_{H}\\ \phi_{L}\end{pmatrix}=\mathcal{O}\begin{pmatrix}a\\ \phi\end{pmatrix}\simeq\begin{pmatrix}a+\epsilon\frac{m_{a,\Lambda}^{2}}{m_{a}^{2}+m_{a,\Lambda}^{2}}\phi\\ \phi-\epsilon\frac{m_{a,\Lambda}^{2}}{m_{a}^{2}+m_{a,\Lambda}^{2}}a\end{pmatrix}\$. (32)
@@ IV.2 Dynamics of the string-wall network
Once $\phi$ begins to evolve due to $V_{\Lambda}$, both $a$ and $\phi$ roll down to the degenerate minima within each domain, and no potential bias remains. In contrast, while $\phi$ is still frozen at its initial value due to its small mass, the potential energy generally varies across different domains. This energy difference acts as a (transient) potential bias, which can induce the collapse of the string-wall network. In this sense, the setup effectively introduces a transient explicit breaking of the PQ symmetry in the early Universe, which disappears later due to the dynamics of the lighter ALP [48].
@@
$N_{\rm DW}=2,\quad n_{a}=1,\quad n_{\phi}=1\ ,$ (33)
in the following analysis, but our results can be straightforwardly applied to more general cases. We further assume that the lighter degree of freedom remains frozen until well after the domain wall network collapses. [^10] In fact, this is a rather conservative condition for the collapse of domain walls. The transient bias could induce a population bias in the domain wall network, which results in a subsequent decay of the domain walls. The decay of domain walls due to population bias for various types of initial fluctuations was studied in Refs. [13, 72]. The onset of $\phi$ oscillation is estimated by
@@
With $V_{\Lambda}$ neglected, domain walls whose tension is determined by $V_{\mathrm{QCD}}$ separate domains where $a\simeq 0$ or $\pi f_{a}$. The domain wall tension is given by
@@
Therefore, the string-wall network decays before the transient bias disappears, if $\epsilon$ is sufficiently small. We then obtain
@@
where we have used $m_{L}\simeq\Lambda^{2}/f_{\phi}$ which follows from $m_{a,\Lambda}\ll m_{a}$. This inequality, together with the assumption $m_{a}>H_{\rm dec}>m_{L}$, can be satisfied when $f_{\phi}\gg f_{a}$.
@@
Requiring that this accounts for the observed dark matter abundance gives
@@
For the QCD axion with a relatively low decay constant ($f_{a}<10^{10}$ GeV) to constitute the dominant dark matter, we must have $T_{\rm dec}<T_{\rm QCD}$. For simplicity, we do not consider a lighter QCD axion ($f_{a}>10^{10}$ GeV), where both the string network and misalignment contributions should be taken into account.
The upper and lower bounds of the ALP decay constant, $f_{\phi,\,\rm max}$ and $f_{\phi,\,\rm min}$, required for successful QCD axion dark matter production are determined by avoiding the overproduction of the light ALP and and by requiring that the domain walls decay due to the transient bias, Eq. (35). At the time of collapse, the axion mass is given approximately by its vacuum value:
$m_{a}(T{{}_{\rm dec}})\approx m_{a0}=2\times\frac{(75.6\,{\rm MeV})^{2}}{f_{a}},$ (43)
and then $\Lambda$ is related to the Hubble parameter at the domain wall network collapse through the condition (38) as
@@
The condition for the transient bias to remain until the domain wall decay, $m_{L}<H_{\rm dec}$, requires
$f_{\phi}>f_{\phi,\,\rm min}\sim 2\times 10^{15}\,{\rm GeV}\left(\frac{1\,{\rm MeV}}{T_{\rm dec}}\right)\left(\frac{10.6}{g_{*,\,\rm dec}}\right)^{1/4}\left(\frac{f_{a}}{10^{8}\,{\rm GeV}}\right)^{1/2}c^{-1/2}\ .$ (45)
The requirement that the light ALP not be overproduced sets the upper limit $f_{\phi,,\rm max}$. The misalignment mechanism produces the lighter ALP, with an abundance given by
@@
Avoiding the overproduction for a typical initial condition $\phi_{\mathrm{ini}}\sim f_{\phi}$ imposes an upper bound on $f_{\phi}$, which corresponds to the largest value of $f_{\phi}$ that yields $\Omega_{L}h^{2}=0.12$. Fig. 5 shows the parameter region where the above scenario works. The blue shaded region corresponds to the overproduction of the light ALP with $f_{a}=2\times 10^{8}\,{\rm GeV}$ for the initial condition $\phi_{\rm ini}=f_{\phi}$. The orange shaded region indicates where the light ALP $A_{L}$ settles to the potential minimum before the domain wall decay. The blue and orange dashed lines show the corresponding cases for $f_{a}=2\times 10^{9}\,{\rm GeV}$. Note that the lower bound on $f_{\phi}$ is a conservative one, since the population bias remaining even after the transient bias is turned off could still drive the domain walls to collapse. With a mild tuning of $\phi_{\rm ini}/f_{\phi}$, one can also suppress $\Omega_{L}$ while keeping $m_{L}$ sufficiently small for the transient bias to be effective. In fact, even when $\phi_{\rm ini}=0$, the transient bias remains nonzero (with $c=1$), as can be seen in Eq. (37). With these caveats in mind, we can see from this figure that the parameter region where the QCD axion constitutes the majority of dark matter is rather small, being limited to $f_{\phi}\approx(2\text{\,--\,}3)\times 10^{15}$ GeV and $T_{\rm dec}\approx 1$ MeV (see Eq. (42)). For higher $T_{\rm dec}$ along the blue solid line, the ALP is the dominant dark matter while the QCD axion is subdominant. The situation does not change for a larger $f_{a}$. This result is expected due to the condition (39). In this scenario, the ALP abundance tends to exceed that of the QCD axion unless the initial displacement $\phi_{\rm ini}$ is fine-tuned.
Lastly, let us mention the isocurvature constraint on $\phi$. Since we assume the pre-inflationary scenario for $\phi$, it acquires quantum fluctuations during inflation, which result in the isocurvature perturbation of $\phi$. This places an upper limit on the inflationary Hubble parameter, but it can be relaxed compared to the one for the QCD axion dark matter scenario since the required decay constant $f_{\phi}$ is larger.
Figure 5: Constraints in the $(f_{\phi},T_{\rm dec})$ plane with $c=0.5$ and $\phi_{\rm ini}=f_{\phi}$. The blue-shaded region indicates overproduction of the light ALP, while the orange-shaded region shows where $A_{L}$ settles to its potential minimum before the domain-wall decay. The solid and dashed lines correspond to $f_{a}=2\times 10^{8}\,{\rm GeV}$ and $2\times 10^{9}\,{\rm GeV}$, respectively.
@@ V Summary and discussion
In this paper, we have proposed two simple solutions to the domain-wall problem induced by the QCD axion by introducing a second, massless (or extremely light) axion:
@@
    String-bundle confinement: If the new axion’s PQ symmetry breaks after inflation, the associated strings can be connected to those of the QCD axion by domain walls. The domain walls then terminate on (and are confined by) these strings, forming cosmologically benign “string bundles,” instead of extended domain walls. These bundles behave as cosmic strings and are much less problematic than domain walls, thereby solving the domain wall problem. Such string bundles can survive to the present epoch and generate cosmic birefringence. If the decay constant $f_{\phi}$ is sufficiently large, the string bundles may also leave observable imprints on the CMB anisotropies as well as stochastic gravitational waves, providing an additional avenue for detection. A simultaneous detection of the QCD axion and a birefringence signal would be a striking experimental signature.
@@
    Transient bias mechanism: If the new axion starts with a homogeneous initial value and is sufficiently light, its potential temporarily biases the QCD-axion potential before the light axion begins to oscillate. This induced bias destabilizes and eliminates the walls while leaving the PQ solution to the strong-CP problem intact. In this case, both the QCD axion and ALP can contribute to dark matter depending on the model parameters.
@@
Related to this, we briefly comment on the connection with the clockwork mechanism. The clockwork mechanism was first applied to the QCD axion in Ref. [39]. In the clockwork QCD axion model, by appropriately assigning charges to many axions, one can generate a large hierarchy between the actual symmetry-breaking scale and the decay constant of the QCD axion. In the minimal setup, the QCD axion is the lightest among the axions, while the others are heavy. Since the clockwork mechanism involves a large number of axions, the internal structure of the corresponding string bundle becomes complicated. In particular, a hierarchical difference is required in the number of constituent axion strings forming the bundle. As a result, such string bundles do not form in realistic cosmological evolution; instead, domain walls associated with the heavy axions are expected to appear [40, 42]. Even in a two-axion system, the formation of string bundles is not guaranteed, and their possible configurations have been classified in Ref. [49]. In this paper, based on such a two-axion framework, we consider a different setup in which the QCD axion is the heavy axion while the lighter one plays the role of an ALP, and we show that in this case the domain-wall problem of the QCD axion can be resolved through the formation of string bundles. We also note that, as mentioned in the main text, a similar model and dynamics in which the QCD axion and the ALP strings form a string bundle have already been pointed out in the context of cosmic birefringence in Ref. [64].
We can also consider the transient bias in a setup where both of the two axions have cosmic strings, and string bundles are not effectively formed due to the domain wall numbers larger than unity. In this case, the light axion becomes highly inhomogeneous, and the bias potential is also inhomogeneous even in a single domain. Thus, it is unclear whether the potential bias can still trigger the collapse of the network. To clarify the dynamics of such systems, dedicated lattice simulations will be necessary, which is left for future work.
@@ Appendix A Typical size of string bundles
as we analytically expected in Sec. III. In the simulation, we choose $f_{\phi}=f_{a}\simeq 7.1m_{A0}$ and constant $m_{A}=m_{A0}$ to synchronize the string formations of both axions and reduce the computational cost.
Figure 6: The snapshot of the string-wall system that appears in the $N_{\rm DW}=2$ case at an early time (left) and a late time (right). The blue(red) triangles indicate the aa-($\phi$-)string locations without distinction of the direction, and the black solid lines indicate the domain wall locations. For reference, we show the Hubble horizon scale using the green square. For simplicity, two kinds of axions have the same decay constant, and the mass is constant in the simulations. One can observe that the string bundles are formed in the late time snapshot.
@@ Appendix B Particle production from topological defects
Fig. 7 shows the time evolution of the heavy axion energy density obtained in lattice simulations. Here, to avoid contamination of the spectra due to the defects, we masked the strings and domain walls. The basic setup is shared with the simulation in Fig. 6, but we choose $f_{\phi}=1.5f_{a}\simeq 10.6m_{A0}$ here. The simulation implies that the final abundance is $2$ – $3$ times larger than the rough analytic estimate shown as the red horizontal dashed line. Thus, our analytic estimate gives the correct axion abundance with an accuracy of ${\cal O}(1)$ factor. We also confirmed that marginally non-relativistic heavy axions are produced at early times, and the comoving energy density freezes at late times.
Let us here comment on possible caveats of the numerical simulations. Due to the radial oscillations of the complex scalar field and the masking method, the abundance exhibits oscillatory behavior at early times, making it difficult to identify when particle production terminates. Nevertheless, we can clearly see that the abundance freezes shortly after $H=H_{\mathrm{bf}}$. Note however that the timing when the string bundles begin to form is relatively close to the onset of the axion oscillations due to computational limitations, and we need more dedicated simulations for quantitative estimates of the final axion abundance. If the axion abundance is enhanced, the decay constant $f_{\phi}$ required to account for the entire dark matter can be smaller, which in turn relaxes the constraints from gravitational signals.
Figure 7: The time evolution of the heavy axion energy density multiplied by the scale factor cubed. The horizontal axis is the Hubble parameter inverse normalized by the heavy axion mass. The black vertical dashed line corresponds to $H=H_{\mathrm{bf}}=\sigma_{A}/\mu_{\phi}$ and the red horizontal dashed line corresponds to the expected abundance assuming instant termination of the particle production at $H=H_{\mathrm{bf}}$. We choose $f_{\phi}=1.5f_{a}\simeq 10.6m_{A0}$. The heavy axion production is deactivated, and the yield freezes at a late time. The oscillations at the early times are considered to be due to the zero mode oscillations in the radial component of the complex scalar, which remain after the masking.